OpenStudy (anonymous):

In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March?

OpenStudy (texaschic101):

x = money in account in march x + 210 = 1.15x 210 = 1.15x - x 210 = 0.15x 210/0.15 = x 1400 = x check.. x + 210 = 1.15x 1400 + 210 = 1.15(1400) 1610 = 1610 (correct) In March, they both had $ 1400. In April, they both have $ 1610